# Unveiling Stripe-shaped Charge Modulations in Doped Mott Insulators

### v1

@@ Abstract
Inspired by recent experimental findings, we investigate various scenarios of the doped Hubbard model with impurity potentials. We calculate the lattice Green’s function in a finite-size cluster and then map it to the continuum real space, which allows for a direct comparison with scanning-tunneling-microscopy measurements on the local density of states. Our simulations successfully reproduce experimental data, including the characteristic stripe- and ladder-shaped structures observed in cuprate systems. Moreover, our results establish a connection between previous numerical findings on stripe-ordered ground states and experimental observations, thereby providing new insights into doped Mott insulators.
@@
Although phase fluctuation theory [1, 3] provides a framework for understanding the phase diagram of cuprates, the microscopic mechanisms driving the transition from Mott insulator to superconductor are not fully understood. A recent scanning tunneling microscopy (STM) experiment [4] on underdoped Bi2201 probes this critical region and observes modulations in local density of states (LDOS) near the Fermi energy. In the insulating sample with a hole concentration of $p=0.08$, there is a spatial phase separation [5, 6] between the hole-free antiferromagnetic phase and the hole-rich phase. The spatial variations within the hole-rich region are highly analogous to those observed in the superconducting sample with $p=0.11$. Specifically, the doped holes self-organize in short-range checkerboard order [7] with a wavelength of approximately $4a_{0}$ (where $a_{0}$ is the lattice constant of the $\rm{CuO_{2}}$ plane), and each plaquette in the checkerboard displays an internal stripe pattern [8, 9]. The primary difference between samples lies in the spatial occupation of the plaquettes. These findings suggest that plaquettes with internal stripe-shaped patterns are essential elementary building blocks. Moreover, even finer spatial structures including ladder- and clover-shaped patterns are observed at higher energy in lightly doped NCCOC [10], which are attributed to molecular orbitals embedded in a Mott insulator [11]. However, systematic investigations are necessary to corroborate their intuitive understanding.
From a theoretical perspective, the Hubbard model and $t$-$J$ model are considered minimal models that qualitatively capture the essence of cuprates [12, 13, 14, 15, 16, 17, 18, 5, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 6]. At low doping levels, the ground state exhibits various charge orders [5, 31, 23], including phase separation, uniform, and stripe configurations with energy close to each other and thus competitive. On the other hand, inhomogeneous impurity potentials of dopants can readily destroy the uniform state in realistic materials, thereby favoring phase separation. To further explain the spatial structures observed in experiments [10, 4, 7], previous studies investigate the ground state of these models in a finite-size cluster. For example, simulations using the density matrix renormalization group (DMRG) on the $t$-$J$ model [17] show that external potential fields can induce charge density modulations in the stripe state [32, 31, 33, 20], resembling the checkerboard state. With the Gutzwiller approximation, the Hubbard model also produces a checkerboard-like pattern [18]. However, differential conductance and LDOS measured with STM are related to the retarded Green’s function in the continuum real space according to the linear response theory [34], which is not directly comparable to those numerically simulated ground state properties such as charge and spin distributions in a lattice system.
@@
[…]
@@
Based on these observations, it is proposed that the clover-shaped pattern reflects the Wannier function of a single-hole state. For example, calculations in density functional theory (DFT) [38, 39, 40] show that the Wannier function exposed to the STM tip exhibits a consistent four-lobe-clover shape for different cuprate materials, regardless of the specific details near the $\rm{CuO_{2}}$ plane. To approximate this Wannier function, we use the symmetrized combination of $p$-orbital wave functions [41]. This choice is consistent with previous research on the sign of wave functions [38, 39, 40], and does not induce qualitative differences in our final results compared to the $d$-orbital wave functions. The density map of this approximated Wannier function $w_{0}(x,y)$ is shown in Fig. 1(c) with $d_{x^{2}-y^{2}}$ symmetry, whose density distribution $|w_{0}(x,y)|^{2}$ is further illustrated in Fig. 1(d), displaying a four-lobe clover shape with its size adjusted to compare with experiments. Following the molecular orbital proposal in doped Mott insulators [10], we consider a linear combination of Wannier functions at different sites in analogy to the linear combination of atomic orbitals (LCAO) [42, 43], with coefficients shown in Fig. 1(e). The resulting molecular orbital has a stripe-shaped distribution (Fig. 1(f)) similar to that observed in experiments.
@@
$H=$ $-t\sum_{\langle ij\rangle,\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}-t^{\prime}\sum_{\langle\langle ij\rangle\rangle,\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}-\mu\sum_{i}n_{i}$ (1)
@@
where we set typical parameters $t=1.0$, $t^{\prime}=-0.3$, and $U=12$ [44, 45, 18, 21]. The number of holes is controlled using $U(1)$ symmetry [46, 47]. The chemical potential $\mu$ is tuned to ensure the ground state with the desired hole number being the true ground state of the entire system [19, 48, 49, 50]. Local potential fields $V_{i}$ are applied to certain lattice sites to simulate impurity effects.
@@
Although the results presented in this manuscript are calculated using a finite cluster with open boundary conditions, we also use a hybrid approach [52] combining CheMPS [37, 35, 36] and cluster perturbation theory (CPT) [53, 54, 55] to study larger systems. The main conclusions remain qualitatively unchanged, suggesting that the physical phenomena observed in the finite cluster are robust and not artifacts of the finite-size simulation.
@@
Numerical results.— We perform simulations of doped holes in finite clusters, including three scenarios based on the number of doped holes: the one-hole, two-hole and four-hole cases. In the one-hole case, we first apply a strong potential field to a single lattice site to pin the doped hole for a $4\times 4$ cluster as shown in Fig. 2(a), whose results are illustrated in Fig. 2(b-d). The density of states (DOS) is shown in Fig. 2(b), where the lower Hubbard band lies below energy 00 and the upper Hubbard band lies above energy $7t_{0}$. Doping-induced in-gap states appear between these two bands, with two distinct peaks marked by red and blue arrows. The real space structures of these peaks are visualized in Fig. 2(c) and Fig. 2(d), respectively. The LDOS for the lower energy peak exhibits a four-lobe clover pattern identical to the shape of the Wannier function, centered at the Cu site where the potential field is applied (Fig. 2(c)). This result indicates that a sufficiently localized hole indeed manifests the shape of the Wannier function, consistent with the experimental results in Fig. 2(e). In contrast, the LDOS for the higher energy peak shows a halved clover pattern in Fig. 2(d), reflecting the transfer of spectral weight from the upper Hubbard band upon doping [56, 57]. These features are qualitatively comparable with Fig. 2(j), while the doped hole spreads more extensively at higher energy in the experimental results, particularly within and outside the clover-shaped pattern.
@@
Moving on to the two-hole case, we aim to verify that the observed stripe-shaped pattern reflects the molecular orbital embedded in the Mott insulator. We consider a $5\times 4$ cluster shown in Fig. 3(a), with impurity potential fields $V_{i}$ applied to the inner $3\times 2$ region and dominant on one side, mimicking the effects of other impurities. The DOS in Fig. 3(b) shows a gap structure near the Fermi energy, with the part below 00 related to the lower Hubbard band and the part above 00 associated with the in-gap state. The peak at energy $0.32t_{0}$ as a characteristic lower-energy state and a higher-energy state at $0.8t_{0}$ for reference are marked by red and blue arrows. The lower-energy LDOS in Fig. 3(c) shows a three-stripe pattern, resembling an alternative LCAO outcome depicted in Fig. 3(e, f), where the Wannier functions form an in-phase superposition analogous to that of a bonding molecular orbital. For the higher-energy state in Fig. 3(d), we see a ladder-shaped distribution of LDOS perpendicular to the lower-energy stripe-shaped distribution, which corresponds to an antibonding molecular orbital and agrees with the experimental results.
@@
We successfully reproduce the stripe-ladder pattern in the two-hole case (more results in Fig. S2 [51]). However, to demonstrate the checkerboard structure with an internal stripe-ladder pattern, we need to consider a minimal four-hole case. In this scenario, a larger number of dopants will induce smoother local potential fields. To simulate this, we trap four holes inside a $10\times 4$ cluster with broad potential fields, as shown in Fig. 4(a). We first calculate the hole densities $1-\langle(n_{i\uparrow}+n_{i\downarrow})\rangle$ and the magnetic moments $\frac{1}{2}\langle(n_{i\uparrow}-n_{i\downarrow})\rangle$ for our ground state in Fig. 4(b). The hole density is concentrated mainly in the central $8\times 2$ region with slight modulations, and this region also forms a domain wall of the surrounding antiferromagnetism [60]. It indicates that the ground state is indeed a stripe state, consistent with previous numerical results for the $t$-$J$ model [17].
@@
Conclusion and discussion.— In summary, we employ the CheMPS method to simulate the retarded lattice Green’s functions for various hole-doping cases in the doped Hubbard model with impurity potential fields. Our calculations enable the projection of the LDOS onto the continuum real space, allowing for a direct and accurate comparison with the STM experiments. We find that the single-hole state exhibits a four-lobe clover pattern in its lowest in-gap state. In contrast, the two-hole state forms a stripe-ladder pattern, which is strikingly similar to the bonding and antibonding molecular orbitals formed by the dimerization of two holes. This result is consistent with experimental observations. Notably, the four-hole state shows two distinct LDOS patterns depending on the energy range. At lower energy, we observe two stripe-shaped plaquettes that are characteristic of the interaction and correlation between the pair of holes. At higher energy, a ladder-shaped structure emerges, in agreement with previous experimental studies [10]. Our calculations also reveal the hole densities and magnetic moments of the ground state, which exhibit a slightly modulated stripe order. This result is consistent with previous numerical calculations using the $t-J$ model [17], further confirming the validity of our simulations.
@@ Acknowledgements.
Acknowledgement.— We thank Shusen Ye and Yayu Wang for helpful discussions and providing STM data. We also thank Zheng-Yu Weng for useful suggestions. This work is supported by the National Natural Science Foundation of China (NSFC) (Grant No. 12174214 and No. 92065205) and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302100).
@@
This Supplemental Material offers an in-depth explanation of the CheMPS method in Appendix A. Furthermore, we conduct a thorough analysis of LDOS distributions under various conditions. Appendices B-D showcase how the real-space patterns in LDOS evolve with these parameters.
Appendix A Chebyshev Matrix Product State Method
@@
Here, $|\psi_{0}\rangle$ is the ground state with energy $E_{0}$. The retarded Green’s function can be obtained by setting $z=\omega+i\eta$ in $G_{ij,\sigma}(z)$. We implement the CheMPS method [35, 36, 37] to calculate the lattice Green’s function. In the following, we illustrate how to compute $G_{ij,\sigma}^{+}(z)$. First, we choose an appropriate energy window $[-\omega_{1},\omega_{2}]$ and perform a rescaling transformation for the variable $z$ and the Hamiltonian $H$.
@@
Appendix B More Results for the one-hole case
@@
Appendix C More Results for the two-hole case
@@
Appendix D More Results for the four-hole case
We conclude our investigation by exploring the behavior of four holes within the lattice. Our analysis begins with the scenario where there are no impurity potential fields, as depicted in Fig. S3(a-e). In this case, the ground state exhibits a uniform hole distribution without magnetic moment. However, the lower-energy LDOS (Fig. S3(d)) still displays two distinct plaquettes. Upon closer inspection, we find that the hole distribution within these plaquettes does not show a striped pattern. As shown in Fig. S3(f), the LDOS distribution along the red cut line in Fig. S3(d) exhibits four peaks with intervals of $1.13a_{0}$, $0.84a_{0}$, and $1.20a_{0}$ from left to right. The side peaks are taller than the middle ones. When a uniform potential field is introduced (Fig. S3(g-l)), we observe that the stripe pattern becomes evident and the stripe order emerges in the ground state. This indicates that the impurity potential can stabilize the stripe order and the corresponding LDOS pattern. As depicted in Fig. S3(l), the LDOS distribution along the red cut line in Fig. S3(j) still shows four peaks, but now the middle two are the tallest, with peak spacings of $1.20a_{0}$. Finally, we modulate the potential field to be stronger on one side (Fig. S3(m-r)), resulting in only three stripes being evident in Fig. S3(r). The rightmost peak is significantly lower than the others, with distances between neighboring peaks ranging from $1.2a_{0}$ to $1.3a_{0}$, consistent with experimental observations. These results further emphasize the significance of impurity potentials and correlation effects between multiple holes.

### v2

# Unveiling Stripe-shaped Charge Density Modulations in Doped Mott Insulators

@@ Abstract
Inspired by recent experimental findings, we investigate various scenarios of the doped Hubbard model with impurity potentials. We calculate the lattice Green’s function in a finite-size cluster and then map it to the continuum real space, which allows for a direct comparison with scanning-tunneling-microscopy measurements on the local density of states. Our simulations successfully reproduce experimental data, including the characteristic stripe- and ladder-shaped structures observed in cuprate systems. Moreover, our results establish a connection between previous numerical findings on stripe-ordered ground states and experimental observations, thus providing new insights into microscopic mechanisms of the Mott insulator to superconductor transition in cuprates.
@@
Although phase fluctuation theory [1, 3] provides a framework for understanding the phase diagram of cuprates, the microscopic mechanisms driving the transition from Mott insulator to superconductor are not fully understood. A recent scanning tunneling microscopy (STM) experiment [4] on underdoped Bi2201 probes this critical region and observes modulations in local density of states (LDOS) near the Fermi energy. In the insulating sample with a hole concentration of $p=0.08$, there is a spatial phase separation [5, 6] between the hole-free antiferromagnetic phase and the hole-rich phase. The spatial variations within the hole-rich region are highly analogous to those observed in the superconducting sample with $p=0.11$. Specifically, the doped holes self-organize in short-range checkerboard order [7, 8] with a wavelength of approximately $4a_{0}$ (where $a_{0}$ is the lattice constant of the $\rm{CuO_{2}}$ plane), and each plaquette in the checkerboard displays an internal stripe pattern [9, 10]. The primary difference between samples lies in the spatial occupation of the plaquettes. These findings suggest that plaquettes with internal stripe-shaped patterns are essential elementary building blocks. Moreover, even finer spatial structures including ladder- and clover-shaped patterns are observed at higher energy in lightly doped NCCOC [11], which are attributed to molecular orbitals embedded in a Mott insulator [12]. However, systematic investigations are necessary to corroborate their intuitive understanding.
From a theoretical perspective, the Hubbard model and $t$-$J$ model are considered minimal models that qualitatively capture the essence of cuprates [13, 14, 15, 16, 17, 18, 19, 5, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 6]. At low doping levels, the ground state exhibits various charge orders [5, 34, 24], including phase separation, uniform, and stripe configurations with energy close to each other and thus competitive. On the other hand, inhomogeneous impurity potentials of dopants can readily destroy the uniform state in realistic materials, thereby favoring phase separation. To further explain the spatial structures observed in experiments [11, 4, 7], previous studies investigate the ground state of these models in a finite-size cluster. For example, simulations using the density matrix renormalization group (DMRG) on the $t$-$J$ model [18] show that external potential fields can induce charge density modulations in the stripe state [35, 34, 36, 21], resembling the checkerboard state. With the Gutzwiller approximation, the Hubbard model also produces a checkerboard-like pattern [19]. However, differential conductance and LDOS measured with STM are related to the retarded Green’s function in the continuum real space according to the linear response theory [37], which is not directly comparable to those numerically simulated ground state properties such as charge and spin distributions in a lattice system.
@@
Figure 1: (Color Online) (a-b) Experimental results of LDOS at $50\textrm{meV}$ (a) and $400\textrm{meV}$ (b), respectively, taken from Ref. [11]. (c-d) Density maps of the approximated Wannier function $w_{0}(x,y)$ (c) and $|w_{0}(x,y)|^{2}$ (d) at the surface exposed to the STM tip. Yellow dots represent Cu sites. (e) The coefficients for LCAO. (f) The density distribution of the LCAO wavefunction.
@@
Based on these observations, it is proposed that the clover-shaped pattern reflects the Wannier function of a single-hole state. For example, calculations in density functional theory (DFT) [41, 42, 43] show that the planar structure of the Wannier function exposed to the STM tip exhibits a consistent four-lobe-clover shape for different cuprate materials, regardless of the specific details near the $\rm{CuO_{2}}$ plane. Moreover, the clover extends over neighboring $\rm{Cu}$ sites due to the spacing between the STM tip and the sample surface [43], in agreement with experimental observations [11]. It also exhibits similarities to localized Zhang-Rice singlets, whose matrix element effect in STM tunneling leads to a four-lobe-clover LDOS distribution [44, 45, 46]. To approximate this four-lobe pattern, we use the symmetrized combination of $p$-orbital wave functions [44]. This choice is consistent with previous research on the sign structure of wave functions [41, 42, 43], and does not induce qualitative differences in our final results compared to the $d$-orbital wave functions. The density map of this approximated Wannier function $w_{0}(x,y)$ is shown in Fig. 1(c) with $d_{x^{2}-y^{2}}$ symmetry, whose density distribution $|w_{0}(x,y)|^{2}$ is further illustrated in Fig. 1(d), displaying a four-lobe clover shape with its size adjusted to compare with experiments. Following the molecular orbital proposal in doped Mott insulators [11], we consider a linear combination of Wannier functions at different sites in analogy to the linear combination of atomic orbitals (LCAO) [47, 48], with coefficients shown in Fig. 1(e). The resulting molecular orbital has a stripe-shaped distribution (Fig. 1(f)) similar to that observed in experiments.
@@
$H=$ $-t_{0}\sum_{\langle ij\rangle,\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}-t_{1}\sum_{\langle\langle ij\rangle\rangle,\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}-\mu\sum_{i}n_{i}$ (1)
@@
where we set typical parameters $t_{0}=1.0$, $t_{1}=-0.3$, and $U=12$ [49, 50, 19, 22]. The number of holes is controlled using $U(1)$ symmetry [51, 52]. The chemical potential $\mu$ is tuned to ensure the ground state with the desired hole number being the true ground state of the entire system [20, 53, 54, 55]. Local potential fields $V_{i}$ are applied to certain lattice sites to simulate impurity effects.
@@
Although the results presented in this manuscript are calculated using a finite cluster with open boundary conditions, we also use a hybrid approach [57] combining CheMPS [40, 38, 39] and cluster perturbation theory (CPT) [58, 59, 60, 61, 62] to study larger systems (see Supplemental Material [56]). The main conclusions remain qualitatively unchanged, suggesting that the physical phenomena observed in the finite cluster are robust and not artifacts of the finite-size simulation.
@@
Numerical results.— We perform simulations of doped holes in finite clusters, including three scenarios based on the number of doped holes: the one-hole, two-hole and four-hole cases. In the one-hole case, we first apply a strong potential field to a single lattice site to pin the doped hole for a $4\times 4$ cluster as shown in Fig. 2(a), whose results are illustrated in Fig. 2(b-d). The density of states (DOS) is shown in Fig. 2(b), where the lower Hubbard band lies below energy 0 and the upper Hubbard band lies above energy $7t_{0}$. Doping-induced in-gap states appear between these two bands, with two distinct peaks marked by red and blue arrows. These in-gap peaks reflect the spectral weight transfer phenomena [63, 64, 65, 66, 67, 8, 68]. In the localized limit, doping a hole leaves behind an empty site where an electron can be added in two possible configurations (spin-up and spin-down). In our model, the added electron must overcome the impurity potential field while interacting with the surrounding antiferromagnetic background, leading to the splitting into two peaks [56]. The real space structures of these peaks are visualized in Fig. 2(c) and Fig. 2(d), respectively. The LDOS for the lower energy peak exhibits a four-lobe clover pattern almost identical to the shape of the Wannier function, centered at the Cu site where the potential field is applied (Fig. 2(c)), consistent with the experimental results in Fig. 2(e). In contrast, the LDOS for the higher energy peak shows a halved clover pattern in Fig. 2(d). These features are qualitatively comparable with Fig. 2(j), while the doped hole spreads more extensively at higher energy in the experimental results, particularly within and outside the clover-shaped pattern.
@@
Moving on to the two-hole case, we aim to verify that the observed stripe-shaped pattern reflects the molecular orbital embedded in the Mott insulator. We consider a $5\times 4$ cluster shown in Fig. 3(a), with impurity potential fields $V_{i}$ applied to the inner $3\times 2$ region and dominant on one side, mimicking the effects of other impurities. The DOS in Fig. 3(b) shows a gap structure near the Fermi energy, with the part below 0 related to the lower Hubbard band and the part above 0 associated with the in-gap state. The peak at energy $0.32t_{0}$ as a characteristic lower-energy state and a higher-energy state at $0.8t_{0}$ for reference are marked by red and blue arrows. The lower-energy LDOS in Fig. 3(c) shows a three-stripe pattern, resembling an alternative LCAO outcome depicted in Fig. 3(e, f), where the Wannier functions form an in-phase superposition analogous to that of a bonding molecular orbital. For the higher-energy state in Fig. 3(d), we see a ladder-shaped distribution of LDOS perpendicular to the lower-energy stripe-shaped distribution, which corresponds to an antibonding molecular orbital and agrees with the experimental results.
@@
We successfully reproduce the stripe-ladder pattern in the two-hole case (more results in Fig. S2 [56]). However, to demonstrate the checkerboard structure with an internal stripe-ladder pattern, we need to consider a minimal four-hole case. In this scenario, a larger number of dopants will induce smoother local potential fields. To simulate this, we trap four holes inside a $10\times 4$ cluster with broad potential fields, as shown in Fig. 4(a). We first calculate the hole densities $1-\langle(n_{i\uparrow}+n_{i\downarrow})\rangle$ and the magnetic moments $\frac{1}{2}\langle(n_{i\uparrow}-n_{i\downarrow})\rangle$ for our ground state in Fig. 4(b). The hole density is concentrated mainly in the central $8\times 2$ region with slight modulations, and this region also forms a domain wall of the surrounding antiferromagnetism [70, 71, 72, 73]. It indicates that the ground state is indeed a stripe state, consistent with previous numerical results for the $t$-$J$ model [18].
@@
Conclusion and discussion.— In summary, we employ the CheMPS method to simulate the retarded lattice Green’s functions for various hole-doping cases in the doped Hubbard model with impurity potential fields. Our calculations enable the projection of the LDOS onto the continuum real space, allowing for a direct and accurate comparison with the STM experiments. We find that the single-hole state exhibits a four-lobe clover pattern in its lowest in-gap state. In contrast, the two-hole state forms a stripe-ladder pattern, which is strikingly similar to the bonding and antibonding molecular orbitals formed by the dimerization of two holes. This result is consistent with experimental observations. Notably, the four-hole state shows two distinct LDOS patterns depending on the energy range. At lower energy, we observe two stripe-shaped plaquettes that are characteristic of the interaction and correlation between the pair of holes. At higher energy, a ladder-shaped structure emerges, in agreement with previous experimental studies [11]. Our calculations also reveal the hole densities and magnetic moments of the ground state, which exhibit a slightly modulated stripe order. This result is consistent with previous numerical calculations using the tJt-J model [18], further confirming the validity of our simulations.
@@ Acknowledgements.
Acknowledgement.— We thank Shusen Ye and Yayu Wang for helpful discussions and providing STM data. We also thank Zheng-Yu Weng for useful suggestions. This work is supported by the National Natural Science Foundation of China (NSFC) (Grant No. 12174214 and No. 12475022) and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302100).
@@
This Supplemental Material provides details about the CheMPS method, LCAO results with different coefficients, LDOS distributions under various conditions, analysis of in-gap states in the one-hole doped case, and results of the hybrid CPT+CheMPS approach.
Appendix A A. Chebyshev Matrix Product State Method
@@
Here, $|\psi_{0}\rangle$ is the ground state with energy $E_{0}$. The retarded Green’s function can be obtained by setting $z=\omega+i\eta$ in $G_{ij,\sigma}(z)$. We implement the CheMPS method [38, 39, 40] to calculate the lattice Green’s function. In the following, we illustrate how to compute $G_{ij,\sigma}^{+}(z)$. First, we choose an appropriate energy window $\omega_{1},\omega_{2}]$ and perform a rescaling transformation for the variable $z$ and the Hamiltonian $H$.
@@
Appendix B B. More results of LCAO
We present two different LCAO results in Fig. 1 and Fig. 3 of the main text, which closely resemble both experimental observations and our many-body numerical simulations. The guiding principles for selecting these LCAO coefficients are based on the symmetries of Wannier functions and molecular orbitals [47, 48]. Within this framework, the most critical factor is the sign structure of the LCAO coefficients, whereas their specific values are less important. This appendix presents various alternative sets of LCAO coefficients along with their corresponding patterns.
First, the in-phase superposition of three atomic orbitals leads to a three-stripe structure, as shown in Fig. S1(a1-a2). However, the side stripes are shorter than the central one. In Fig. S1(b1-b2), we introduce six additional atomic orbitals on the sides, which possess opposite signs relative to the central three. This modification results in an extension of the side stripes. A similar three-stripe pattern is observed in Fig. S1(c1-c2) and Fig. S1(h1-h2). Even when the LCAO coefficients are highly irregular but maintain the same sign structure, as seen in Fig. S1(d1-d2), the resulting pattern remains consistent with experimental features. In contrast, when the difference in sign is removed, the side stripes become rounded and the ends of the central stripe appear broadened (Fig. S1(e1-e2)). Furthermore, when the correct sign structure is not maintained, the three-stripe structure is completely replaced by a larger clover pattern, as observed in Fig. S1(f1-f2) and Fig. S1(g1-g2). These findings underscore the importance of maintaining the correct sign structure in the LCAO coefficients to faithfully reproduce the symmetry properties of the molecular orbitals and achieve consistency with the experimental results.
Figure S1: Different choices of LCAO coefficients and the corresponding molecular orbital density distributions.
Appendix C C. More Results for the one-hole case
@@
Appendix D D. More Results for the two-hole case
@@
Appendix E E. More Results for the four-hole case
We conclude our investigation by exploring the behavior of four holes within the lattice. Our analysis begins with the scenario where there are no impurity potential fields, as depicted in Fig. S4(a-e). In this case, the ground state exhibits a uniform hole distribution without a magnetic moment. However, the lower-energy LDOS (Fig. S4(d)) still displays two distinct plaquettes. Upon closer inspection, we find that the hole distribution within these plaquettes does not show a striped pattern. As shown in Fig. S4(f), the LDOS distribution along the red cut line in Fig. S4(d) exhibits four peaks with intervals of $1.13a_{0}$, $0.84a_{0}$, and $1.20a_{0}$ from left to right. The side peaks are taller than the middle ones. When a uniform potential field is introduced (Fig. S4(g-l)), we observe that the stripe pattern becomes evident and the stripe order emerges in the ground state. This indicates that the impurity potential can stabilize the stripe order and the corresponding LDOS pattern. As depicted in Fig. S4(l), the LDOS distribution along the red cut line in Fig. S4(j) still shows four peaks, but now the middle two are the tallest, with peak spacings of $1.20a_{0}$. Finally, we modulate the potential field to be stronger on one side (Fig. S4(m-r)), resulting in only three stripes being evident in Fig. S4(r). The rightmost peak is significantly lower than the others, with distances between neighboring peaks ranging from $1.2a_{0}$ to $1.3a_{0}$, consistent with experimental observations. These results further emphasize the significance of impurity potentials and correlation effects between multiple holes.